# Dependability of UAV-Based Networks and Computing Systems: A Survey

### v1

@@ 1. Introduction
Dependability has long been a cornerstone of cyber-physical systems (Miclea and Sanislav, 2011). Traditional CPS frameworks emphasize fault tolerance, redundancy, and rigorous verification to mitigate failures. UAV-based systems, as a specialized class of CPS, inherit these principles but face unique challenges due to their aerial mobility, limited energy resources, and exposure to dynamic environmental conditions (e.g., wind gusts, electromagnetic interference) (Pacheco et al., 2021a; Xue et al., 2023b; Zhan and Xu, 2024). For instance, while ground-based CPS may rely on stable power sources and fixed communication infrastructure, UAVs must balance limited battery life with computational demands, all while maintaining connectivity in ad-hoc airborne networks. These fundamental differences require a re-examination of existing dependability paradigms to effectively address the specific vulnerabilities and operational demands of UAV-based systems.
Dependability in UAV-based systems encompasses a range of attributes, including reliability, availability, safety, and resilience, all of which are essential to maintaining Quality of Service (QoS) and accomplishing mission objectives (Gupta and Gupta, 2022; Farrukh and West, 2023; Avizienis et al., 2004). These systems operate in environments where they face diverse sources of uncertainty, such as hardware and software failures, network disruptions, energy constraints, environmental disturbances, and communication breakdowns. For instance, a UAV performing a search-and-rescue mission may encounter battery depletion or signal loss, potentially leading to mission failure or safety hazards (Kaymaz et al., 2024). Recent studies have identified energy constraints as a primary threat, as battery depletion often causes mission aborts in long-duration operations (Lin et al., 2021a). Morereover, network disruptions in UAV swarms can hinder coordination and data sharing due to intermittent connectivity (Alioua et al., 2018).
@@ 5. Dependability threats
Energy Constraint | Battery Degradation | (Lu et al., 2022), (Hou et al., 2020c), (Yang et al., 2022), (Li and Nie, 2023) |
@@
Data Packet Loss | (Mukherjee et al., 2023), (Wang et al., 2019c), (Nwadiugwu et al., 2021), (Jiang et al., 2019), (Haber et al., 2019a), (Bai et al., 2023b), (Zhang et al., 2023d), (Chu et al., 2021a), (Mukherjee et al., 2020) | |
@@
Network Congestion | (Picano and Fantacci, 2024a), (Nwadiugwu et al., 2021), (Picano and Fantacci, 2024a), (She and Ouyang, 2021) | |
@@
Hardware Failure | Sensor Malfunction | (Saadaoui et al., 2023), (Hou et al., 2020b), (Shen et al., 2023b), (Ju et al., 2024a), (Huang et al., 2024a), (Yuan et al., 2022), (Chen et al., 2020b), (Kabashkin, 2023) |
@@
Actuator Failure | (Amin and Obermaisser, 2018a), (Zhou and Cao, 2018), (Wang et al., 2019c), (Wang et al., 2021f), (Hirota et al., 2020a), (Hu and Seiler, 2015) | |
@@
Component Aging | (Shao et al., 2023a), (Aloqaily et al., 2021), (Haber et al., 2019b), (Dogea et al., 2023), (Ren et al., 2017), (Guo et al., 2022a), (Zheng et al., 2023), (Wang et al., 2020c), (D’Agati et al., 2024), (Moeyersons et al., 2021), (Xiong et al., 2019), (Li et al., 2019), (Navardi and Mohsenin, 2023a), (Zhang, 2020), (Jiang and He, 2023), (Zheng et al., 2021), (Faraci et al., 2023a), (Liu et al., 2024d), (Seisa et al., 2022a), (Luo et al., 2021a), (Haber et al., 2019a), (Bao et al., 2023), (Jiang et al., 2019), (Falcão et al., 2023a), (Bai et al., 2023c), (Jung et al., 2023a), (Ngoenriang et al., 2022a), (Feng et al., 2024), (Geng et al., 2024a), (Ning et al., 2024), (De Paula Soares et al., 2023), (Ren et al., 2024), (He et al., 2023a), (Gupta and Gupta, 2022), (Chen et al., 2024a), (Picano and Fantacci, 2024b), (Wang et al., 2021a),(Li et al., 2023b), (Feng et al., 2024), (Khoramnejad et al., 2024), (Zhou et al., 2024b), (Lei et al., 2023), (Shao et al., 2023a), (Hou et al., 2020d), (Yu et al., 2023a), (Zhang et al., 2023d), (Tripathi et al., 2023a), (Lorincz et al., 2021a), (Pacheco et al., 2021a), (Patnayak et al., 2021a), (Pamuklu et al., 2023), (Kaleem et al., 2019a), (Zeng et al., 2024b), (Adil et al., 2024), (Hang et al., 2020a), (Wang et al., 2022c), (Li et al., 2024a) | |
@@
Mechanical Breakdown | (Luo et al., 2021b), (Zhang et al., 2021a), (Russell et al., 2018), (Shen et al., 2023b), (Lv et al., 2024), (Wang et al., 2020b), (Dai et al., 2024), (Choutri et al., 2022), (Zhou et al., 2024a), (Lindqvist et al., 2020), (Kabashkin, 2023), (Lorincz et al., 2021b), (Behjat et al., 2019) | |
@@
Software Failure | System Vulnerabilities | (Pradhan et al., 2015), (Wang et al., 2023a), (Ma et al., 2024), (Vamvakas et al., 2019), (Wang et al., 2023a), (Ren et al., 2017), (Venkatagiri et al., 2018a), (Qu et al., 2022), (Chang et al., 2024), (Watanabe and Machida, 2022), (Zhong et al., 2021), (Zuo et al., 2022), (Zhou et al., 2023), (Chen et al., 2022a), (Yang et al., 2019), (Hirota et al., 2020a) |
@@
System Crashes | (Ma et al., 2024), (Amin and Obermaisser, 2018a), (Hirota et al., 2020a), (Kramer and Stubberud, 2021), (Liu et al., 2019), (Qu et al., 2022), (Chang et al., 2024), (Pan et al., 2024), (Jennings et al., 2022), (Shang et al., 2023), (Taleb et al., 2023a), (Zhang et al., 2021c), (Grasso et al., 2022b), (Ritchie et al., 2020), (Lin et al., 2021b), (Mo et al., 2022), (Gupta et al., 2021), (Hong et al., 2022), (Zhou et al., 2024d), (Zhong et al., 2021), (Nguyen et al., 2023), (Zhang and Li, 2018), (Kaymaz et al., 2024), (Ahamad and Mishra, 2024), (Goel et al., 2021), (Bai et al., 2023b) | |
@@
Environmental Impact | Adverse Weather Conditions | (Wang et al., 2021c), (Wang et al., 2021d), (Gallego-Madrid et al., 2020), (Hossain et al., 2017), (Pradhan et al., 2015), (Haber et al., 2021a), (Yang et al., 2022), (Lee et al., 2022a), (Sankarasrinivasan et al., 2015), (Barnawi et al., 2021), (Mcilwaine and Rivas Casado, 2021), (Yang et al., 2022), (Fan et al., 2023), (Venkatagiri et al., 2018b), (Kumar and Jannesari, 2023) |
@@
Terrain Challenges | (Dai et al., 2018), (Rokaha et al., 2019) , (Paranjothi et al., 2019), (Bakhtiari et al., 2020a), (Gao et al., 2022), (Peng et al., 2023), (Dai et al., 2018), (Kim et al., 2024b), (Faraci et al., 2023a), (Mao et al., 2023), (Shang et al., 2023), (Lv et al., 2024), (Wang et al., 2021f), (Hadi and Ghazizadeh, 2023), (Zhou et al., 2024a) | |
@@
Operational Failures | Human Error | (Xu and Zhai, 2020), (Hou et al., 2020b), (Yu et al., 2021), (Liu et al., 2023c), (Yuan et al., 2022), (Falcão et al., 2023a), (Kim et al., 2024b), (Mao et al., 2023), (Faraci et al., 2023b), (Thapliyal et al., 2024a), (Zhang et al., 2024d), (Xu and Zhai, 2020), (Manzoor et al., 2021), (Jeong et al., 2020a), (He et al., 2023b), (Hui et al., 2024), (Falcão et al., 2023a), (Li et al., 2024e), (Kaleem et al., 2019b), (Ozger et al., 2018a), (Rao and Naidu, 2024), (Ouyang et al., 2023), (Ju et al., 2024a), (Tang et al., 2023), (Amos et al., 2021), (Shen et al., 2023a), (Wang et al., 2020a), (Vamvakas et al., 2019), (Asim et al., 2022) |
@@
Maintenance Issues | (Guo et al., 2022b), (Jo et al., 2018), (Wu et al., 2024), (Lv et al., 2024), (Yang et al., 2017), (Ren et al., 2017), (Zhang, 2020), (Pradhan et al., 2015), (Lin et al., 2021b), (Zhang and Zhang, 2024), (Patnayak et al., 2021a), (Shams et al., 2023), (Huang and Wang, 2024) |
@@ 5.1. Energy Constraint
Battery Degradation: Battery degradation refers to the gradual loss of battery capacity and performance due to repeated charging and discharging cycles. In UAVs, lithium-polymer (LiPo) batteries are the most commonly used due to their high energy density (Lu et al., 2022). However, over time, these batteries undergo chemical aging, reducing the amount of charge they can hold (Hou et al., 2020c). This results in shorter flight durations (Yang et al., 2022), inconsistent power output, and potential mid-flight failures (Li and Nie, 2023). Factors that accelerate degradation include overcharging, deep discharges, and exposure to high temperatures. For instance, a UAV that initially offers 30 minutes of flight time may reduce to just 15 minutes after 100 charge cycles. Real-time monitoring and battery health management systems can help detect and mitigate degradation risks.
@@ 5.2. Network Issues
Data Packet Loss: Packet loss occurs when one or more transmitted data packets fail to reach their destination, leading to gaps in communication. In UAVs, this can result in missing telemetry, sensor data, or delayed command execution. Causes include RF interference, poor signal quality, and buffer overflow in the UAV or ground station. In high-mobility networks such as FANETs (Flying Ad Hoc Networks), the likelihood of packet loss and intermittent connections increases due to dynamic topology changes and unstable links (Mukherjee et al., 2023). Additionally, quadrotor UAV systems are vulnerable to network issues such as random delays and packet loss, which can destabilize real-time control and disrupt reliable wireless communication (Wang et al., 2019c). Enhancements at the MAC protocol level, such as prioritizing nodes based on proximity, have been proposed to reduce network delay and improve robustness against packet loss (Nwadiugwu et al., 2021). Poor packet scheduling and resource allocation in multi-agent UAV systems can lead to higher packet loss, especially when network conditions change (Jiang et al., 2019). Network disconnections between UAVs, ground stations, satellites, and IoT devices are also identified as major contributors to data loss and increased latency, impacting the reliability of computational offloading and communication services (Haber et al., 2019a; Bai et al., 2023b; Zhang et al., 2023d; Chu et al., 2021a). High packet loss can degrade mission accuracy. In IoDT applications, high message latency and packet loss can reduce quality of service (QoS), especially in dynamic environments (Mukherjee et al., 2020). Limited energy and memory in edge devices make these challenges worse. For example, a UAV sending image data back to a ground station may deliver incomplete or corrupted files, requiring retransmission and causing delays. Protocols with error correction and adaptive retransmission strategies help mitigate these issues.
@@
Network Congestion: Network congestion happens when multiple UAVs or devices compete for the same communication resources, causing delays and data collisions. This is common in swarm operations, public events, or urban areas. Congestion leads to increased latency, packet loss, and sometimes total communication failure (Picano and Fantacci, 2024a). Lack of effective node prioritization in UAV networks can worsen congestion by increasing competition for limited communication channels (Nwadiugwu et al., 2021). Heavy traffic and strict QoS demands on UAV and ground edge nodes can cause congestion, delay communications, and lead to missed deadlines for important tasks (Picano and Fantacci, 2024a; She and Ouyang, 2021). For instance, in a drone light show involving dozens of UAVs, congestion could cause synchronization errors, impacting performance and safety. Solutions involve dynamic channel allocation, time-division multiplexing, and efficient communication protocols that adapt to network load.
@@ 5.3. Hardware failure
Sensor Malfunction: Sensor malfunctions can occur in components such as GPS, IMU, altimeters, or cameras. These devices are critical for navigation, stabilization, and mission-specific tasks. Faulty or drifting sensors provide inaccurate data to the flight controller, leading to instability or incorrect responses (Saadaoui et al., 2023). For example, a GPS module reporting an incorrect position may cause the drone to veer off course, while an altimeter failure could result in improper altitude hold. Environmental factors such as magnetic interference or moisture can also contribute to sensor errors. UAV systems’ sensor malfunctions can disrupt critical services such as mobile edge computing, vehicular communications, or fog computing operations, impacting overall system reliability (Hou et al., 2020b; Shen et al., 2023b; Ju et al., 2024a; Huang et al., 2024a; Yuan et al., 2022; Chen et al., 2020b). Environmental conditions, such as bad weather or collisions, further increase the risk of sensor malfunctions, requiring fault management strategies like deploying additional UAVs to replace malfunctioning ones and maintain service continuity (Kabashkin, 2023). To address these risks, system designs increasingly incorporate redundancy, regular calibration procedures, health monitoring, and fault-tolerant offloading strategies.
@@
Actuator Failure: Actuators, such as servos and motor controllers, translate flight control commands into mechanical motion. Failure in these components can result in reduced maneuverability or total loss of control (Amin and Obermaisser, 2018a). A jammed servo may lock a control surface on a fixed-wing UAV, while a faulty motor controller can shut down a rotor on a quadcopter. Main causes include overheating, mechanical wear, electrical faults, or broader hardware issues affecting actuators during flight operations (Zhou and Cao, 2018; Wang et al., 2019c; Amin and Obermaisser, 2018a). Some systems address actuator failures by enabling rapid re-planning in response to failures, allowing continued operation through dynamic adjustments (Hirota et al., 2020a). Low-cost UAVs are especially vulnerable due to limited actuator redundancy, making accurate fault detection and control switching essential; recent approaches use reliability frameworks with pivotal decomposition and new fault detection and isolation (FDI) metrics to handle actuator failures without relying on extra hardware (Hu and Seiler, 2015). Quadrotor UAV control systems are particularly vulnerable to actuator failures when combined with environmental disturbances and network issues like random delays and packet loss, which can threaten stability and disrupt real-time control (Wang et al., 2021f). Ensuring component quality, regular inspection, and incorporating fail-safes like motor redundancy in quadcopters can mitigate such risks.
@@
Component Aging: With repeated usage, UAV components degrade due to mechanical fatigue (Shao et al., 2023a), corrosion, and thermal cycling. Bearings in motors, solder joints on PCBs, and electrical connectors are particularly vulnerable. Aging components can reduce efficiency, increase energy consumption, or suddenly fail mid-flight. In addition, long-term stress and environmental exposure can gradually weaken critical UAV hardware, increasing the risk of unexpected failures and service disruptions (Aloqaily et al., 2021; Haber et al., 2019b; Dogea et al., 2023; Ren et al., 2017; Guo et al., 2022a; Zheng et al., 2023; Wang et al., 2020c; D’Agati et al., 2024; Moeyersons et al., 2021; Xiong et al., 2019). Such aging-related failures, combined with network disconnections between aircraft and ground stations, may result in increased latency or data loss, impacting the reliability of data transmission and service continuity. Efficient energy management is also critical, as aging components may draw more power, reducing operational time. Moreover, aging systems can be more prone to software faults, where bugs or vulnerabilities emerge over time, requiring robust maintenance and development practices to uphold dependability (Ren et al., 2017).
@@
Mechanical Breakdown: Structural components like frames, propellers, and landing gear are exposed to significant stress from takeoffs, landings, and potential collisions. Mechanical fatigue and wear over time can lead to cracks and fractures, particularly in components like landing gear and propeller mounts (Luo et al., 2021b; Zhang et al., 2021a; Russell et al., 2018). Stress accumulation from repeated missions and vibrations contributes to deformation and weakening of airframes and connectors (Shen et al., 2023b; Lv et al., 2024; Wang et al., 2020b). Mid-air collisions or impacts during landing are common triggers for abrupt structural failures, especially in multi-UAV or urban operations where maneuvering space is limited (Dai et al., 2024; Choutri et al., 2022; Zhou et al., 2024a; Lindqvist et al., 2020). Additionally, poor material quality or inadequate structural design increases the risk of breakage under operational loads (Kabashkin, 2023; Lorincz et al., 2021b; Behjat et al., 2019). A broken propeller mid-flight, for example, can result in immediate loss of control, particularly in systems lacking redundancy. Mechanical breakdowns are often preventable through visual inspections, structural stress testing, and the use of high-quality materials during design and assembly.

### v2

@@ 1. Introduction
Dependability has long been a cornerstone of cyber-physical systems (Miclea and Sanislav, 2011). Traditional CPS frameworks emphasize fault tolerance, redundancy, and rigorous verification to mitigate failures. UAV-based systems, as a specialized class of CPS, inherit these principles but face unique challenges due to their aerial mobility, limited energy resources, and exposure to dynamic environmental conditions (e.g., wind gusts, electromagnetic interference) (Pacheco et al., 2021; Xue et al., 2023b; Zhan and Xu, 2024). For instance, while ground-based CPS may rely on stable power sources and fixed communication infrastructure, UAVs must balance limited battery life with computational demands, all while maintaining connectivity in ad-hoc airborne networks. These fundamental differences require a re-examination of existing dependability paradigms to effectively address the specific vulnerabilities and operational demands of UAV-based systems.
Dependability in UAV-based systems encompasses a range of attributes, including reliability, availability, safety, and resilience, all of which are essential to maintaining Quality of Service (QoS) and accomplishing mission objectives (Gupta and Gupta, 2022; Farrukh and West, 2023; Avizienis et al., 2004). These systems operate in environments where they face diverse sources of uncertainty, such as hardware and software failures, network disruptions, energy constraints, environmental disturbances, and communication breakdowns. For instance, a UAV performing a search-and-rescue mission may encounter battery depletion or signal loss, potentially leading to mission failure or safety hazards (Kaymaz et al., 2024). Recent studies have identified energy constraints as a primary threat, as battery depletion often causes mission aborts in long-duration operations (Lin et al., 2021). Morereover, network disruptions in UAV swarms can hinder coordination and data sharing due to intermittent connectivity (Alioua et al., 2018).
@@ 5. Dependability threats
| Energy Constraint | Battery Degradation | (Lu et al., 2022), (Hou et al., 2020b), (Yang et al., 2022), (Li and Nie, 2023) |
@@
| Data Packet Loss | (Mukherjee et al., 2023), (Wang et al., 2019b), (Nwadiugwu et al., 2021), (Jiang et al., 2019), (Haber et al., 2019), (Bai et al., 2023a), (Zhang et al., 2023c), (Chu et al., 2021), (Mukherjee et al., 2020) | |
@@
| Network Congestion | (Picano and Fantacci, 2024), (Nwadiugwu et al., 2021), (Picano and Fantacci, 2024), (She and Ouyang, 2021) | |
@@
| Hardware Failure | Sensor Malfunction | (Saadaoui et al., 2023), (Hou et al., 2020a), (Shen et al., 2023b), (Ju et al., 2024), (Huang et al., 2024a), (Yuan et al., 2022), (Chen et al., 2020b), (Kabashkin, 2023) |
@@
| Actuator Failure | (Amin and Obermaisser, 2018), (Zhou and Cao, 2018), (Wang et al., 2019b), (Wang et al., 2021d), (Hirota et al., 2020), (Hu and Seiler, 2015) | |
@@
| Component Aging | (Shao et al., 2023), (Aloqaily et al., 2021), (Haber et al., 2019), (Dogea et al., 2023), (Ren et al., 2017), (Guo et al., 2022a), (Zheng et al., 2023), (Wang et al., 2020c), (D’Agati et al., 2024), (Moeyersons et al., 2021), (Xiong et al., 2019), (Li et al., 2019), (Navardi and Mohsenin, 2023), (Zhang, 2020), (Jiang and He, 2023), (Zheng et al., 2021), (Faraci et al., 2023), (Liu et al., 2024c), (Seisa et al., 2022), (Luo et al., 2021), (Haber et al., 2019), (Bao et al., 2023), (Jiang et al., 2019), (Falcão et al., 2023), (Bai et al., 2023b), (Jung et al., 2023a), (Ngoenriang et al., 2022), (Feng et al., 2024), (Geng et al., 2024), (Ning et al., 2024), (De Paula Soares et al., 2023), (Ren et al., 2024), (He et al., 2023), (Gupta and Gupta, 2022), (Chen et al., 2024a), (Picano and Fantacci, 2024), (Wang et al., 2021a),(Li et al., 2023b), (Feng et al., 2024), (Khoramnejad et al., 2024), (Zhou et al., 2024b), (Lei et al., 2023), (Shao et al., 2023), (Hou et al., 2020b), (Yu et al., 2023a), (Zhang et al., 2023c), (Tripathi et al., 2023), (Lorincz et al., 2021), (Pacheco et al., 2021), (Patnayak et al., 2021), (Pamuklu et al., 2023), (Kaleem et al., 2019), (Zeng et al., 2024b), (Adil et al., 2024), (Hang et al., 2020), (Wang et al., 2022c), (Li et al., 2024a) | |
@@
| Mechanical Breakdown | (Luo et al., 2021), (Zhang et al., 2021a), (Russell et al., 2018), (Shen et al., 2023b), (Lv et al., 2024), (Wang et al., 2020b), (Dai et al., 2024), (Choutri et al., 2022), (Zhou et al., 2024a), (Lindqvist et al., 2020), (Kabashkin, 2023), (Lorincz et al., 2021), (Behjat et al., 2019) | |
@@
| Software Failure | System Vulnerabilities | (Pradhan et al., 2015), (Wang et al., 2023a), (Ma et al., 2024), (Vamvakas et al., 2019), (Wang et al., 2023a), (Ren et al., 2017), (Venkatagiri et al., 2018), (Qu et al., 2022), (Chang et al., 2024), (Watanabe and Machida, 2022), (Zhong et al., 2021), (Zuo et al., 2022), (Zhou et al., 2023), (Chen et al., 2022a), (Yang et al., 2019), (Hirota et al., 2020) |
@@
| System Crashes | (Ma et al., 2024), (Amin and Obermaisser, 2018), (Hirota et al., 2020), (Kramer and Stubberud, 2021), (Liu et al., 2019), (Qu et al., 2022), (Chang et al., 2024), (Pan et al., 2024), (Jennings et al., 2022), (Shang et al., 2023), (Taleb et al., 2023), (Zhang et al., 2021c), (Grasso et al., 2022b), (Ritchie et al., 2020), (Lin et al., 2021), (Mo et al., 2022), (Gupta et al., 2021), (Hong et al., 2022), (Zhou et al., 2024c), (Zhong et al., 2021), (Nguyen et al., 2023), (Zhang and Li, 2018), (Kaymaz et al., 2024), (Ahamad and Mishra, 2024), (Goel et al., 2021), (Bai et al., 2023a) | |
@@
| Environmental Impact | Adverse Weather Conditions | (Wang et al., 2021b), (Wang et al., 2021b), (Gallego-Madrid et al., 2020), (Hossain et al., 2017), (Pradhan et al., 2015), (Haber et al., 2021), (Yang et al., 2022), (Lee et al., 2022a), (Sankarasrinivasan et al., 2015), (Barnawi et al., 2021), (Mcilwaine and Rivas Casado, 2021), (Yang et al., 2022), (Fan et al., 2023), (Venkatagiri et al., 2018), (Kumar and Jannesari, 2023) |
@@
| Terrain Challenges | (Dai et al., 2018), (Rokaha et al., 2019) , (Paranjothi et al., 2019), (Bakhtiari et al., 2020), (Gao et al., 2022), (Peng et al., 2023), (Dai et al., 2018), (Kim et al., 2024b), (Faraci et al., 2023), (Mao et al., 2023), (Shang et al., 2023), (Lv et al., 2024), (Wang et al., 2021d), (Hadi and Ghazizadeh, 2023), (Zhou et al., 2024a) | |
@@
| Operational Failures | Human Error | (Xu and Zhai, 2020), (Hou et al., 2020a), (Yu et al., 2021), (Liu et al., 2023c), (Yuan et al., 2022), (Falcão et al., 2023), (Kim et al., 2024b), (Mao et al., 2023), (Faraci et al., 2023), (Thapliyal et al., 2024), (Zhang et al., 2024d), (Xu and Zhai, 2020), (Manzoor et al., 2021), (Jeong et al., 2020), (He et al., 2023), (Hui et al., 2024), (Falcão et al., 2023), (Li et al., 2024e), (Kaleem et al., 2019), (Ozger et al., 2018), (Rao and Naidu, 2024), (Ouyang et al., 2023), (Ju et al., 2024), (Tang et al., 2023), (Amos et al., 2021), (Shen et al., 2023a), (Wang et al., 2020a), (Vamvakas et al., 2019), (Asim et al., 2022) |
@@
| Maintenance Issues | (Guo et al., 2022b), (Jo et al., 2018), (Wu et al., 2024), (Lv et al., 2024), (Yang et al., 2017), (Ren et al., 2017), (Zhang, 2020), (Pradhan et al., 2015), (Lin et al., 2021), (Zhang and Zhang, 2024), (Patnayak et al., 2021), (Shams et al., 2023), (Huang and Wang, 2024) |
@@ 5.1. Energy Constraint
Battery Degradation: Battery degradation refers to the gradual loss of battery capacity and performance due to repeated charging and discharging cycles. In UAVs, lithium-polymer (LiPo) batteries are the most commonly used due to their high energy density (Lu et al., 2022). However, over time, these batteries undergo chemical aging, reducing the amount of charge they can hold (Hou et al., 2020b). This results in shorter flight durations (Yang et al., 2022), inconsistent power output, and potential mid-flight failures (Li and Nie, 2023). Factors that accelerate degradation include overcharging, deep discharges, and exposure to high temperatures. For instance, a UAV that initially offers 30 minutes of flight time may reduce to just 15 minutes after 100 charge cycles. Real-time monitoring and battery health management systems can help detect and mitigate degradation risks.
@@ 5.2. Network Issues
Data Packet Loss: Packet loss occurs when one or more transmitted data packets fail to reach their destination, leading to gaps in communication. In UAVs, this can result in missing telemetry, sensor data, or delayed command execution. Causes include RF interference, poor signal quality, and buffer overflow in the UAV or ground station. In high-mobility networks such as FANETs (Flying Ad Hoc Networks), the likelihood of packet loss and intermittent connections increases due to dynamic topology changes and unstable links (Mukherjee et al., 2023). Additionally, quadrotor UAV systems are vulnerable to network issues such as random delays and packet loss, which can destabilize real-time control and disrupt reliable wireless communication (Wang et al., 2019b). Enhancements at the MAC protocol level, such as prioritizing nodes based on proximity, have been proposed to reduce network delay and improve robustness against packet loss (Nwadiugwu et al., 2021). Poor packet scheduling and resource allocation in multi-agent UAV systems can lead to higher packet loss, especially when network conditions change (Jiang et al., 2019). Network disconnections between UAVs, ground stations, satellites, and IoT devices are also identified as major contributors to data loss and increased latency, impacting the reliability of computational offloading and communication services (Haber et al., 2019; Bai et al., 2023a; Zhang et al., 2023c; Chu et al., 2021). High packet loss can degrade mission accuracy. In IoDT applications, high message latency and packet loss can reduce quality of service (QoS), especially in dynamic environments (Mukherjee et al., 2020). Limited energy and memory in edge devices make these challenges worse. For example, a UAV sending image data back to a ground station may deliver incomplete or corrupted files, requiring retransmission and causing delays. Protocols with error correction and adaptive retransmission strategies help mitigate these issues.
@@
Network Congestion: Network congestion happens when multiple UAVs or devices compete for the same communication resources, causing delays and data collisions. This is common in swarm operations, public events, or urban areas. Congestion leads to increased latency, packet loss, and sometimes total communication failure (Picano and Fantacci, 2024). Lack of effective node prioritization in UAV networks can worsen congestion by increasing competition for limited communication channels (Nwadiugwu et al., 2021). Heavy traffic and strict QoS demands on UAV and ground edge nodes can cause congestion, delay communications, and lead to missed deadlines for important tasks (Picano and Fantacci, 2024; She and Ouyang, 2021). For instance, in a drone light show involving dozens of UAVs, congestion could cause synchronization errors, impacting performance and safety. Solutions involve dynamic channel allocation, time-division multiplexing, and efficient communication protocols that adapt to network load.
@@ 5.3. Hardware failure
Sensor Malfunction: Sensor malfunctions can occur in components such as GPS, IMU, altimeters, or cameras. These devices are critical for navigation, stabilization, and mission-specific tasks. Faulty or drifting sensors provide inaccurate data to the flight controller, leading to instability or incorrect responses (Saadaoui et al., 2023). For example, a GPS module reporting an incorrect position may cause the drone to veer off course, while an altimeter failure could result in improper altitude hold. Environmental factors such as magnetic interference or moisture can also contribute to sensor errors. UAV systems’ sensor malfunctions can disrupt critical services such as mobile edge computing, vehicular communications, or fog computing operations, impacting overall system reliability (Hou et al., 2020a; Shen et al., 2023b; Ju et al., 2024; Huang et al., 2024a; Yuan et al., 2022; Chen et al., 2020b). Environmental conditions, such as bad weather or collisions, further increase the risk of sensor malfunctions, requiring fault management strategies like deploying additional UAVs to replace malfunctioning ones and maintain service continuity (Kabashkin, 2023). To address these risks, system designs increasingly incorporate redundancy, regular calibration procedures, health monitoring, and fault-tolerant offloading strategies.
@@
Actuator Failure: Actuators, such as servos and motor controllers, translate flight control commands into mechanical motion. Failure in these components can result in reduced maneuverability or total loss of control (Amin and Obermaisser, 2018). A jammed servo may lock a control surface on a fixed-wing UAV, while a faulty motor controller can shut down a rotor on a quadcopter. Main causes include overheating, mechanical wear, electrical faults, or broader hardware issues affecting actuators during flight operations (Zhou and Cao, 2018; Wang et al., 2019b; Amin and Obermaisser, 2018). Some systems address actuator failures by enabling rapid re-planning in response to failures, allowing continued operation through dynamic adjustments (Hirota et al., 2020). Low-cost UAVs are especially vulnerable due to limited actuator redundancy, making accurate fault detection and control switching essential; recent approaches use reliability frameworks with pivotal decomposition and new fault detection and isolation (FDI) metrics to handle actuator failures without relying on extra hardware (Hu and Seiler, 2015). Quadrotor UAV control systems are particularly vulnerable to actuator failures when combined with environmental disturbances and network issues like random delays and packet loss, which can threaten stability and disrupt real-time control (Wang et al., 2021d). Ensuring component quality, regular inspection, and incorporating fail-safes like motor redundancy in quadcopters can mitigate such risks.
@@
Component Aging: With repeated usage, UAV components degrade due to mechanical fatigue (Shao et al., 2023), corrosion, and thermal cycling. Bearings in motors, solder joints on PCBs, and electrical connectors are particularly vulnerable. Aging components can reduce efficiency, increase energy consumption, or suddenly fail mid-flight. In addition, long-term stress and environmental exposure can gradually weaken critical UAV hardware, increasing the risk of unexpected failures and service disruptions (Aloqaily et al., 2021; Haber et al., 2019; Dogea et al., 2023; Ren et al., 2017; Guo et al., 2022a; Zheng et al., 2023; Wang et al., 2020c; D’Agati et al., 2024; Moeyersons et al., 2021; Xiong et al., 2019). Such aging-related failures, combined with network disconnections between aircraft and ground stations, may result in increased latency or data loss, impacting the reliability of data transmission and service continuity. Efficient energy management is also critical, as aging components may draw more power, reducing operational time. Moreover, aging systems can be more prone to software faults, where bugs or vulnerabilities emerge over time, requiring robust maintenance and development practices to uphold dependability (Ren et al., 2017).
@@
Mechanical Breakdown: Structural components like frames, propellers, and landing gear are exposed to significant stress from takeoffs, landings, and potential collisions. Mechanical fatigue and wear over time can lead to cracks and fractures, particularly in components like landing gear and propeller mounts (Luo et al., 2021; Zhang et al., 2021a; Russell et al., 2018). Stress accumulation from repeated missions and vibrations contributes to deformation and weakening of airframes and connectors (Shen et al., 2023b; Lv et al., 2024; Wang et al., 2020b). Mid-air collisions or impacts during landing are common triggers for abrupt structural failures, especially in multi-UAV or urban operations where maneuvering space is limited (Dai et al., 2024; Choutri et al., 2022; Zhou et al., 2024a; Lindqvist et al., 2020). Additionally, poor material quality or inadequate structural design increases the risk of breakage under operational loads (Kabashkin, 2023; Lorincz et al., 2021; Behjat et al., 2019). A broken propeller mid-flight, for example, can result in immediate loss of control, particularly in systems lacking redundancy. Mechanical breakdowns are often preventable through visual inspections, structural stress testing, and the use of high-quality materials during design and assembly.